How do you multiply a rational expression like this one: 25-x^2/12*6/5-x? Whats the answer?: How do you multiply a rational expression like this one: 25-x^2/12*6/5-x? Whats the answer?
Answer by oscargut(667)

[25-x^2/12]=
[6(25-x^2)/12(5-x)]=
(25-x^2)/2(5-x)=(5-x)(5+x)/2(5-x)=(5+x)/2
Answer: (x+5)/2